OpenStudy (anonymous):

using linearization formula, get an approximation for cos(91°), with f(x)=cos(x), a=90° (in radians) and x=91° (in radians)

OpenStudy (anonymous):

If y = cos x, x in radians, then dy = -sin x dx. For x = pi/2 and dx = pi/180. dy = -sin(pi/2)(pi/180) = -pi/180. So cos(91 degrees) = cos(pi/2) - pi/180 = -pi/180.